How do i get the F(16) From this problem
F(x)= 3x-9 if x <(or equal to)-2
-2/5x if x > -2

Answer by rothauserc(4718)

F(x)= 3x-9 if x <(or equal to)-2
-2/5x if x > -2
note that 16 > -2
we use -2/5x = -2/5*16 = -1/40 = -0.025
F(16) = -0.025